# The DSA Transparency Database: Auditing Self-reported Moderation Actions by Social Media

### v2

@@ 4.3. Content Type
The required attribute content type is a list with one or more of the following predefined values: app, audio, image, product, synthetic media (e.g., content generated via AI), text, video, and other. Despite the possibility to indicate multiple values, all eight platforms only specified a single value for all of their SoRs, except YouTube, which however indicated multiple values only in a handful of cases ($<$0.01%). We found that social media platforms greatly differ in how they specify the content type of their SoRs, as shown on the left side of Figure 4. Pinterest and LinkedIn only set content type other, and did not map any of their internal types to one or more of the predefined values in the DSA-TDB. YouTube also had a remarkable share of other content (93.9%), prevalently ads. Facebook and Instagram have similar distributions, with most content being other (more than 55%), all detailed as accounts, which is consistent with their distributions of decision types in Table 2. Incidentally, Facebook was the only social media with content type product (1.8%). More than half of Snapchat’s moderated content was video, with other content (24%) being mainly accounts. Still, 16.6% of other was labelled as multi-media, instead of using multiple predefined values as prescribed by the DSA-TDB. Interestingly, TikTok’s moderated content was mostly text (55.5%) and not video (33.5%), with a small share of other content (1.3%), primarily accounts. Unexpectedly, moderated content by X was overwhelmingly synthetic media (99.8%), with the second being audio (0.01%) —the only social media to use this type— and only a handful of text.
Figure 4. Distribution of moderated content type by platform. For the detailed other content type (right side of the figure), labels were simplified for space reasons.
@@ 4.4. Timeliness
Artice 17 of the DSA requires that platforms communicate their SoRs without undue delay to allow close to real-time updates. Pinterest, Snapchat, and TikTok started submitting SoRs on the very day that the DSA-TDB was made available —that is, on 25 September 2023. Facebook, Instagram, YouTube, and LinkedIn the following day. X was last in line and started a week later, on 3 October 2023. As sketched in Figure 5, we then analyzed the delay with which each social media communicated their SoRs, measured as the difference (in days) between the date of application of the moderation intervention and the date of creation of the corresponding SoR. In addition, a prompt response by platforms in moderating an infringing content is arguably desirable. Hence, we also measured the application delay (in days) of the moderation decision with respect to the creation of the infringing content.
@@
Regarding communication delay, LinkedIn, Pinterest, and Snapchat communicated their SoRs on the same day of application with little to no fluctuations, as shown in Figure 6. Despite the late start, X caught up on the first week of submissions by communicating content moderated since 25 September. Afterwards, it began to communicate SoRs the same day of their application. On the contrary, TikTok started by stably submitting on the same day of application, but in the second half of the 100-day period it communicated SoRs with an uprising mean delay of up to eight days. Facebook and Instagram communicated their SoRs with a mean delay of one day, with a few delay surges common to both Meta platforms. On the other hand, YouTube had a minimum delay of one day, but with marked fluctuations throughout the period of analysis.
Figure 6. Platform-wise daily mean of the delay (measured in days) in the communication of statements of reasons to the DSA-DTB.
Concerning application delay, we observed great heterogeneity in its distribution among platforms. For this reason, in their analysis and visualization we opted for using few categorical cuts, as shown in Figure 7. In addition to visualizing the distributions of applications delay, we also report in the figure the 80% winsorized mean (W) as a robust measure of centrality that lessens the influence of outliers. Oddly, X was the only platform that always applied a decision on the same day the infringing content was created (W$=$0). TikTok had the second lowest delay, with 89% of decisions being applied the day of content creation. Next, LinkedIn applied its decision within 7 days in 89.8% of their SoRs, while Instagram in 81.7%. Snapchat had a similar value for the same timeframe (81.4%), but in 14.1% of SoRs it took more than 30 days to apply its decision. Facebook presented a similar but more marked situation, with 72% and 21.5%, respectively. YouTube, on the other hand, had a substantial amount of decisions that took more than 30 days (70.7%), while also having many within 15–30 days (17.5%). However, the distribution with the highest mean and dispersion was that of Pinterest, as 17.5% of decisions were applied the same day of content creation while 70.8% after 30 days, with many even taking hundreds of days.
@@ 4.5. Automated Means
Automation can be applied at various steps in the content moderation process. It is often used to scale up the detection of offending content and to speed up the decision on how to moderate it. Consequently, the DSA-TDB includes two attributes to describe the possible use of automation for moderation: automated detection (yes, no) and automated decision (not, partially, fully). Figure 8 provides a combined view of the use of automation for detection and decision by social media platforms. Again, we see different approaches being adopted. Surprisingly, none of the SoRs submitted by X specified the use of automated means whatsoever. Snapchat is second in terms of non-automated means in their SoRs (74%). Pinterest mainly uses non-automated detection with partially automated decision (98%), whereas LinkedIn mainly uses automated detection together with non-automated decision (64%). YouTube is the only platform that uses all combinations of automated means, with automated detection and non-automated decision being the most common (51%), followed by fully automated means (28%), and non-automated means (16%). Once more, Instagram and Facebook have similar distributions, with automated detection and partially automated decision being the most common for both, with 93% and 99%, respectively. Almost opposite of X, TikTok predominantly indicates fully automated means for moderation (92%).
@@
To provide a holistic view of social media behavior with respect to automated moderation, we converted the boolean and ordinal attributes of the DSA-TDB into two numeric indices of automated detection and automated decision. Both indices are defined in the [0, 1] range and are valued, for each platform, depending on the frequency with which automated detection and decision occurred in the platform’s SoRs. We were then able to position each platform in the bi-dimensional space depicted in Figure 9, where dot size is proportional to the number of SoRs per MAAR, and dot color represents the mean of the automated detection and automated decision indices. The platform positioning within the newly defined automation space reveals four groups of platforms with overall similar degrees of automation: low (X and Snapchat), modest (Pinterest), substantial (LinkedIn, YouTube, Instagram, and Facebook), and high (TikTok). At the platform level, the ratio in SoRs per MAAR is moderately correlated with an automated index ($r=0.6)$. However, the correlation between number of SoRs per MAAR and automated detection is much lower ($r=0.3$) compared to automated decision ($r=0.9$). In particular, Pinterest stands apart from others platforms due to its very low level of automated detection but very high level of partially automated decision.
Figure 9. Use of automated means by social media platforms for the detection of infringements and decision on these. Size is proportional to the number of statements of reasons (SoRs) per monthly authenticated active recipients (MAAR). In general, the bigger this ratio is, the higher the automated index ($r=0.6)$, with Pinterest being somewhat of an outlier due to its marginal automated detection.
@@ 4.6. Comparison with Transparency Reports
TikTok, YouTube, Snapchat. TikTok’s transparency report lacks several details and their data only covers the month of September 2023, shortcomings that are acknowledged therein. Based on the more detailed actions for removed content, we found a considerable automation discrepancy. The report specifies that 45% of non-ad content was removed automatically, whereas in the DSA-TDB it was 95%, more than double for a partially overlapping period. In the case of YouTube, even if it is the only platform to use all combinations of automated means without one being predominant —in accordance with its transparency report— we found intriguing anomalies. In the DSA-TDB we saw a remarkable spike in SoRs applied on 5 October 2023, reaching 732.4K. For reference, YouTube’s daily median is 112.8K. Almost all of these are ads disabled because deemed outside the scope of the platform, but via non-automated means, which seems implausible. Furthermore, there were three similar spikes in mid December with even higher single-day volumes (¿1.25M each) corresponding to a majority of SoRs having automated detection but non-automated decision. Unfortunately, with the data available we could not delve deeper into these anomalies. Snapchat’s transparency report is somewhat peculiar, as they already had in place a namesake global transparency report, released twice a year. Hence, they only integrated this report via a document with some high-level information specific to the EU, in order to comply with the first mandatory DSA reporting period. Nevertheless, we noticed an odd inconsistency. DSA transparency reports must include data about Trusted Flaggers: legal entities responsible for notifying platforms of illegal content (European Parliament and Council, 2022). As per Article 19(2) of the DSA, Trusted Flaggers are officially appointed by certain EU bodies. However, at the time of writing no Trusted Flagger has yet been designated. For this reason, no platform reported in their SoRs to have received notifications from Trusted Flaggers, with the exception of Snapchat, for which we found a few hundreds SoRs with notification source type set as Trusted Flagger. Upon further inspection we realized that Snapchat has a similar namesake internal program, [^8] https://values.snap.com/safety/safety-center unrelated to the DSA’s one. This homonymy probably caused some confusion in Snapchat’s reporting to the DSA-TDB, where they referred to their internal program instead of the intended one.
@@ 5. Discussion
We first discuss the adequacy of the data submitted to the DSA-TDB (RQ1). We then examine inconsistencies in the submitted data (RQ2) and we describe interesting differences and similarities among platforms (RQ3). Finally, we ponder on the potential misuse of the database contents in its current state, for research purposes and policymaking.
@@ 5.1. RQ1: Adequacy of Submitted Data
We noticed several inadequacies in the data submitted by multiple social media platforms. In several cases the attributes in the database were conceived to allow platforms a higher level of granularity in characterizing a SoR, compared to what is actually done. For instance, there is a severe lack of detail regarding the types of infringements. The optional category addition is practically never used, and the rich set of available keywords in category specification is severely underutilized. Further, in most cases platforms specified the infringement category with the least informative scope of platform service. Similarly, platforms have the possibility to apply more than one of the four types of restriction decisions to a SoR, but none of the social media did so. In particular, it was surprising to see that no monetary restriction decision was specified whatsoever, especially by platforms with well-established content monetization schemes subject to strong moderation, such as YouTube (Ma and Kou, 2022). Likewise, content type allows multiple values, but in the great majority of instances only one was set. This is particularly problematic with synthetic media, as it would be important to also know the kind of media of the generated content —whether text, audio, video, or a combination of them (Alam et al., 2022). These widespread deficiencies not only diminish the utility of the DSA-TDB by limiting the depth of available data, but also compromise its commitment to transparency.
@@
In general, the initial implementation of the SoR communication systems by social media platforms was not as comprehensive as it could and should be. Content language is an illustrating example. The language in which moderators work is one the key aspects that VLOPs must cover in their transparency reports —an expected requirement given the multilingual policy to which the EU strives (Gazzola, 2016). Still, none of the social media platforms set the content language in their SoRs. From a technical point of view, most if not all of these very large social media already have mechanisms to register or automatically detect the language of a piece of content or account (e.g., via the user interface settings). Moreover, this attribute is essential for the platform, not only for moderation via manual our automated means (Lees et al., 2022), but also to sell targeted ads and generate revenue (Beauvisage et al., 2023). We presume platforms avoid communicating this attribute —and the resulting workload— simply because it is optional.
@@
That being said, we believe that some fault also lies in the shortcomings of the current database schema. We acknowledge the challenges entailed in the definition of a common and unifying schema, not only for several social media, but for all the kinds of online platforms covered by the DSA. In fact, in mid 2023 there was a public consultation [^9] https://digital-strategy.ec.europa.eu/en/library/digital-services-act-summary-report-public-consultation-dsa-transparency-database on the DSA-TDB, in which 48 stakeholders participated, including representatives of 6 VLOPs. The consultation was intended to refine an existing draft of the schema and SoR communication mechanisms. This also means that VLOPs only had a few months to implement these changes. Nonetheless, we noticed that certain attribute definitions could be improved, based on their actual use.
For instance, many platforms apply their restriction decisions at the account level, not because of a specific type of multi-media content created by it, but due to the account itself. However, in the initial database schema it was not possible to specify an account as the target of a restriction. Hence, many platforms indicated the SoR content type as other and detailed it as an account-based type, as we showed in Figure 4. Perhaps, for the future it would be better to separate the target kind at least into two attributes: content and account, both of which could then be further specified, as it already occurs for content type and the many account variants in other. On a similar note, instead of being a content type, synthetic media could be a separate boolean attribute synthetic target. As another example, we believe that automated detection should be an ordinal variable with the same three levels as automated decision instead of the current dichotomous attribute. Finally, a moderation action might have been subject to an appeal process, in which case an SoR update would be necessary. Indeed, the DSA-TDB allows updating an SoR in case personal data is mistakenly present. Nonetheless, indicating a (successful or not) appeal process is not currently possible. Appropriately enough, the EU has undertaken an open approach to further develop the DSA-TDB and states it is willing to receive and act upon feedback.

### v4

@@ 4.3. Content Type
The required attribute content type (§2 Specifications of the content affected by the decision) is a list with one or more of the following predefined values: app, audio, image, product, synthetic media (e.g., content generated via AI), text, video, and other. Despite the possibility to indicate multiple values, as also mentioned in the official documentation with multiple examples, almost all platforms only specified a single value for all of their SoRs. The only exception is YouTube, which however indicated multiple values only in a handful of cases ($<$0.01%). We found that social media platforms greatly differ in how they specify the content type of their SoRs, as shown on the left side of Figure 4. Pinterest and LinkedIn only set content type other, and did not map any of their internal types to one or more of the predefined values in the DSA-TDB. YouTube also had a remarkable share of other content (93.9%), prevalently ads. Facebook and Instagram have similar distributions, with most content being other (more than 55%), all detailed as accounts, which is consistent with their distributions of decision types in Table 2. Incidentally, Facebook was the only social media with content type product (1.8%). More than half of Snapchat’s moderated content was video, with other content (24%) being mainly accounts. Still, 16.6% of other was labelled as multi-media, instead of using multiple predefined values as prescribed by the DSA-TDB. Interestingly, TikTok’s moderated content was mostly text (55.5%) and not video (33.5%), with a small share of other content (1.3%), primarily accounts. Unexpectedly, moderated content by X was overwhelmingly synthetic media (99.8%), with the second being audio (0.01%) —the only social media to use this type— and only a handful of text.
[…]
Stacked barchart ordered by distribution of moderated content type. Both Pinterest and LinkedIn have 100% of their moderated content as other, followed by YouTube with 93.9%. Both Facebook and Instagram have slightly more than half of their content type as other. On the other hand, most of Snapchat’s moderated content type is video, whereas for TikTok was less than half, with text being more than half. X is the most dissimilar, with almost all of its moderated content being synthetic media.
Figure 4. Distribution of moderated content type by platform. For the detailed other content type (right side of the figure), we simplified the original free text labels for space reasons.
@@ 4.4. Timeliness
Artice 17 of the DSA requires that platforms communicate their SoRs without undue delay to allow close to real-time updates. Pinterest, Snapchat, and TikTok started submitting SoRs on the very day that the DSA-TDB was made available —that is, on 25 September 2023. Facebook, Instagram, YouTube, and LinkedIn the following day. X was last in line and started a week later, on 3 October 2023. As depicted in Figure 5, we then analyzed the delay with which each social media communicated their SoRs, measured as the difference (in days) between the date of application of the moderation intervention and the date of creation of the corresponding SoR. In addition, a prompt response by platforms in moderating an infringing content is arguably desirable. Hence, we also measured the application delay (in days) of the moderation decision with respect to the creation of the infringing content. To compute the communication and application delays we use the content creation date (§2.3 Date on which the content was created on the online platform), the moderation intervention application date (§4.1 Application date), and the SoR creation date, which is among the available metadata.
[…]
The diagram illustrates three events: creation of infringing content, application of moderated decision, and creation of statement of reasons. The application and communication delays are the respective lags between these three events.
@@
Regarding communication delay, LinkedIn, Pinterest, and Snapchat communicated their SoRs on the same day of application with little to no fluctuations, as shown in Figure 6. Despite the late start, X caught up on the first week of submissions by communicating content moderated since 25 September. Afterwards, it began to communicate SoRs the same day of their application. On the contrary, TikTok started by stably submitting on the same day of application, but in the second half of the 100-day period it communicated SoRs with an uprising mean delay of up to eight days. Facebook and Instagram communicated their SoRs with a mean delay of one day, with a few delay spikes common to both Meta platforms. On the other hand, YouTube had a minimum delay of one day, but with marked fluctuations throughout the period of analysis.
[…]
Faceted time series of daily mean delay by platform for the 100 days of the database. Facebook and Instagram have very similar time series, both presenting a relatively low delay but with a noticeable spike at the end. LinkedIn, Pinterest, and Snapchat have a minimal and steady delay. TikTok started with a minimal delay but halfway started to increase up to almost 8 days by the end. X started with a delay of almost 8 days, but after a few days it stabilized at almost zero. Finally, YouTube had a minimal delay at the start and end, but in the middle there many fluctuations in the delay, reaching up to 4 days.
Figure 6. Platform-wise daily mean of the delay (measured in days) in the communication of statements of reasons to the DSA-TDB.
Concerning application delay, we observed great heterogeneity in its distribution among platforms. For this reason, in their analysis and visualization we opted for using a few categorical cuts, as shown in Figure 7. In addition to visualizing the distributions of applications delay, we also report in the figure the 80% winsorized mean (W) as a robust measure of centrality that lessens the influence of outliers. Oddly, X was the only platform that always applied a decision on the same day the infringing content was created (W$=$0). TikTok had the second lowest delay, with 89% of decisions being applied the day of content creation. Next, LinkedIn applied its decision within 7 days in 89.8% of their SoRs, while Instagram in 81.7%. Snapchat had a similar value for the same timeframe (81.4%), but in 14.1% of SoRs it took more than 30 days to apply its decision. Facebook presented a similar but more marked situation, with 72% and 21.5%, respectively. YouTube, on the other hand, had a substantial amount of decisions that took more than 30 days (70.7%), while also having many within 15–30 days (17.5%). However, the distribution with the highest mean and dispersion was that of Pinterest, as 17.5% of decisions were applied the same day of content creation while 70.8% after 30 days, with many even taking hundreds of days.
[…]
Faceted barcharts of the delay distribution. X is the only platform with a zero-day delay in all of its statements or reasons, with TikTok following closely. LinkedIn, Instagram, Snapchat, and Facebook have a zero-day delay in more than half of their application decisions. On the other hand, for both YouTube and Pinterest most of their delays are of 30 days or more.
@@ 4.5. Automated Means
Automation can be applied at various steps in the content moderation process. It is often used to scale up the detection of offending content and to speed up the decision on how to moderate it. Consequently, the DSA-TDB includes two attributes to describe the possible use of automation for moderation (§9 Automated detection and §10 Automated decision): automated detection (yes, no) and automated decision (not, partially, fully). Figure 8 provides a combined view of the use of automation for detection and decision by social media platforms. Again, we see different approaches being adopted. Surprisingly, none of the SoRs submitted by X specified the use of automated means whatsoever. Snapchat is second in terms of non-automated means in their SoRs (74%). Pinterest mainly uses non-automated detection with partially automated decision (98%), whereas LinkedIn mainly uses automated detection together with non-automated decision (64%). YouTube is the only platform that uses all combinations of automated means, with automated detection and non-automated decision being the most common (51%), followed by fully automated means (28%), and non-automated means (16%). Once more, Instagram and Facebook have similar distributions, with automated detection and partially automated decision being the most common for both, with 93% and 99%, respectively. Almost opposite of X, TikTok predominantly indicates fully automated means for moderation (92%).
[…]
Faceted heatmap by platform, with rows indicating three automation decision levels (full, partial, and not) and columns indication automated detection (no or yes). At one extreme, X is the only platform that has all of its statements of reason as not using any automated means. At the other extreme, TikTok indicates that 92% of its statements of reasons use full automated decision and detection means.
@@
To provide a holistic view of social media behavior with respect to automated moderation, we converted the boolean and ordinal attributes of the DSA-TDB into two numeric indices of automated detection and automated decision. Both indices are defined in the [0, 1] range and are valued, for each platform, depending on the frequency with which automated detection and decision occurred in the platform’s SoRs. We were then able to position each platform in the bi-dimensional space depicted in Figure 9, where dot size is proportional to the number of SoRs per MAAR, and dot color represents the mean of the automated detection and automated decision indices. The platform positioning within the newly defined automation space reveals four groups of platforms with overall similar degrees of automation: low (X and Snapchat), modest (Pinterest), substantial (LinkedIn, YouTube, Instagram, and Facebook), and high (TikTok). At the platform level, the ratio in SoRs per MAAR is moderately correlated with the automation index ($r=0.6)$. However, the correlation between number of SoRs per MAAR and automated detection is much lower ($r=0.3$) compared to automated decision ($r=0.9$). In particular, Pinterest stands apart from others platforms due to its very low level of automated detection but very high level of partially automated decision.
[…]
Scatterplot in which the eight platforms are distributed on a square panel based on their level of automated decision index and automated detection index. X, Snapchat, and Pinterest are on the quadrant for low automated means. LinkedIn, YouTube, Instagram and Facebook are on the quadrant with high detection but low automated detection. TikTok is the only platform present in the quadrant with high automation in terms of detection and decision. Incidentally, X and TikTok are almost diametrically opposite in terms of automated means.
Figure 9. Use of automated means by social media platforms for the detection of infringements and decision on these. Size is proportional to the number of statements of reasons (SoRs) per monthly authenticated active recipients (MAAR). In general, the bigger this ratio is, the higher the automation index ($r=0.6)$, with Pinterest being somewhat of an outlier due to its marginal automated detection.
@@ 4.6. Comparison with Transparency Reports
TikTok, YouTube, Snapchat. TikTok’s transparency report lacks several details and their data only covers the month of September 2023, shortcomings that are acknowledged therein. Based on the more detailed actions for removed content, we found a considerable automation discrepancy. The report specifies that 45% of non-ad content was removed automatically, whereas in the DSA-TDB it was 95%, more than double for a partially overlapping period. In the case of YouTube, even if it is the only platform to use all combinations of automated means without one being predominant —in accordance with its transparency report— we found intriguing anomalies. In the DSA-TDB we saw a remarkable spike in SoRs applied on 5 October 2023, reaching 732.4K. For reference, YouTube’s daily median is 112.8K. Almost all of these are ads disabled because deemed outside the scope of the platform, but via non-automated means, which seems implausible. Furthermore, there were three similar spikes in mid December with even higher single-day volumes (¿1.25M each) corresponding to a majority of SoRs having automated detection but non-automated decision. Unfortunately, with the data available we could not delve deeper into these anomalies. Snapchat’s transparency report is somewhat peculiar, as they already had in place a namesake global transparency report, released twice a year. Hence, they only integrated this report via a document with some high-level information specific to the EU, in order to comply with the first mandatory DSA reporting period. Nevertheless, we noticed an odd inconsistency. DSA transparency reports must include data about Trusted Flaggers: legal entities responsible for notifying platforms of illegal content (European Parliament and Council, 2022). As per Article 19(2) of the DSA, Trusted Flaggers are officially appointed by certain EU bodies. However, at the time the SoRs were submitted to the DSA-TDB, no Trusted Flagger had yet been designated. For this reason, no platform reported in their SoRs to have received notifications from Trusted Flaggers (§7.1 Information source), with the exception of Snapchat, for which we found a few hundreds SoRs with notification source type set as Trusted Flagger. Upon further inspection we realized that Snapchat has a similar namesake internal program, [^9] https://values.snap.com/safety/safety-center unrelated to the DSA’s one. This homonymy probably caused some confusion in Snapchat’s reporting to the DSA-TDB, where they referred to their internal program instead of the intended one. Since then, some Trusted Flaggers have been appointed pursuant Article 19(2) of the DSA. [^10] https://digital-strategy.ec.europa.eu/en/policies/trusted-flaggers-under-dsa Future analyses of the notifications submitted by Trusted Flaggers under the DSA could therefore be meaningful, but should be careful to discard Snapchat’s initial unreliable data.
@@ 5. Discussion
We first discuss the adequacy of the data submitted to the DSA-TDB (RQ1), next we examine inconsistencies in the submitted data (RQ2), and then we describe interesting differences and similarities among platforms (RQ3). Based on our findings concerning these aspects, we make suggestions on how the current structure of the database could be improved. Finally, we describe the main limitations of our study, and ponder on the potential implications of the database and the DSA in general on future online moderation research.
@@ 5.1. RQ1: Adequacy of Submitted Data
We noticed several inadequacies in the data submitted by multiple social media platforms. In several cases the attributes in the database were conceived to allow platforms a higher level of granularity in characterizing a SoR, compared to what is actually done. For instance, there is a severe lack of detail regarding the types of infringements. The optional category addition is practically never used, and the rich set of available keywords in category specification is severely underutilized. Further issues arise because a single moderation action can fall under multiple categories. For example, the removal of sexual content could be categorized under pornography or sexualized content and further specified as either adult sexual material or image-based sexual abuse (§16 Category & Specification in the official documentation). However, in some cases the same content might also be categorized as scope of platform service, which includes a nudity subcategory, or as protection of minors, under child sexual abuse material. Although the documentation encourages platforms to be as specific as possible and to include multiple categories and specifications, this almost never happens. As a result, a certain degree of ambiguity exists in how different platforms chose to categorize the moderated content. This leads to challenges in interpreting DSA-TDB records, as it is currently impossible to distinguish whether two moderation actions were categorized differently due to actual differences in the moderated content or due to inconsistencies in how platforms reported their actions. In particular, the category scope of platform service has a higher level of uncertainty due to its multiple unrelated subcategories. From the analysis of Figure 2, we note however that all platforms either indicated pornography or sexualized content or scope of platform service as the most frequent type of infringement. The presence of the nudity subcategory in scope of platform service could possibly explain this result, indicating a general and platform-independent tendency to prevalently moderate adult content. However, due to the missing optional fields that further specify infringement categories, particularly those more wide-ranging such as scope of platform service, it is not possible to draw a definitive conclusion in this regard.
@@
Similarly, platforms have the possibility to apply more than one of the four types of restriction decisions to a SoR (§3 The type of restriction(s) imposed), but none of the social media did so. In particular, it was surprising to see that no monetary restriction decision was specified whatsoever, especially by platforms with well-established content monetization schemes subject to strong moderation, such as YouTube (Ma and Kou, 2022). Likewise, content type allows multiple values (§2 Specifications of the content affected by the decision), but in the great majority of instances only one was set. This is particularly problematic with synthetic media, as it would be important to also know the kind of media of the generated content —whether text, audio, video, or a combination of them (Alam et al., 2022). Again, while the official documentation explicitly suggests including as many types as are applicable to the content, this was almost never done. These widespread deficiencies not only diminish the utility of the DSA-TDB by limiting the depth of available data, but also compromise its commitment to transparency.
@@
In general, the initial implementation of the SoR communication systems by social media platforms was not as comprehensive as it could and should be. Content language is an illustrating example (§2.4 Language of the content). The language in which moderators work is one of the key aspects that VLOPs must cover in their transparency reports —an expected requirement given the multilingual policy to which the EU strives (Gazzola, 2016). Still, none of the social media platforms set the content language in their SoRs. From a technical point of view, most if not all of these very large social media already have mechanisms to register or automatically detect the language of a piece of content or account (e.g., via the user interface settings). Moreover, this attribute is essential for the platform, not only for moderation via manual our automated means (Lees et al., 2022), but also to sell targeted ads and generate revenue (Beauvisage et al., 2023). We presume platforms avoid communicating this attribute —and the resulting workload— simply because it is optional.
@@
That being said, we believe that some fault also lies in the shortcomings of the current database schema. We acknowledge the challenges entailed in the definition of a common and unifying schema, not only for several social media, but for all the kinds of online platforms covered by the DSA. In fact, in mid 2023 there was a public consultation [^11] https://digital-strategy.ec.europa.eu/en/library/digital-services-act-summary-report-public-consultation-dsa-transparency-database on the DSA-TDB, in which 48 stakeholders participated, including representatives of 6 VLOPs. The consultation was intended to refine an existing draft of the schema and SoR communication mechanisms. This also means that VLOPs only had a few months to implement these changes. Nonetheless, we noticed that certain attribute definitions could be improved, based on their actual use. For instance, many platforms apply their restriction decisions at the account level, not because of a specific type of multi-media content created by it, but due to the account itself. However, in the initial database schema it is not possible to specify an account as the target of a restriction (§2.1 Type of content affected). Hence, many platforms indicated the SoR content type as other and detailed it with a free text label related to accounts, as we show in Figure 4.
Such inadequacy of the database schema leads to confusion as to what is the actual target of a moderation action. Indeed, since the other content type is detailed in a free text field with all sorts of heterogeneous labels, including various kinds of accounts and groups, we were unable to accurately distinguish among these for further analysis. The issue also propagates to other related database fields, such as the creation date of the infringing target. As indicated in the documentation (§2.3 Date on which the content was created on the online platform), this might be related to “the date that specific content was posted, or the date that a user account was registered” —an attribute of two inherently different entities conflated into a single field. This and other ambiguities can have important repercussions on downstream analyses. For instance, in Figure 7 we show the histograms of the platform-wise distribution of the delays between the application date of a moderation action and the creation date of the corresponding moderated target. With the current DSA-TDB schema it is not possible to confidently assess whether the relatively high values on the extreme histogram bins (zero-day delay and more than 30 days delay) in some platforms (e.g., Facebook, Instagram, Pinterest) is due to the the inherently different creation times among different targets (e.g., content, account, groups), or due to some other phenomena. This practical example underlines the implications of a poorly specified database schema and calls for additional specification effort. Appropriately enough, the EU has undertaken an open approach to further develop the DSA-TDB and states it is willing to receive and act upon feedback.